Mr.sparrow was terrified when he opened his door and saw 88 legs scurrying across the floor. he quickly counted and noticed that there were a total of 12 spiders and cockroaches. how many spider and cockroaches were there.
Answer by fractalier(6550)

Call the number of spiders and cockroaches, s and c. Then we write
8s + 6c = 88
s + c = 12
Now multiply the bottom one by 8 and subtract it...we get
8s + 6c = 88
-(8s + 8c = 96)
----------------
-2c = -8
c = 4 cockroaches
s = 8 spiders